Consider the relationship between the number of bids an item on eBay received and the items selling price. The following is a sample of 5 items sold through an auction. Price in Dollars 29 32 40 45 49 Number of Bids 7 89 Table Copy Data Step 3 of 5: Calculate the estimated variance of slope, s. Round your answer to three decimal places. Answer(How to Enter) 2 Points Keypad BTables Next
Step 4 of 5: Construct the 95 % confidence interval for the slope. Round your answers to three decimal places. Answer(How to Enter) 4 Points Keypad Tables Lower endpoint: Next Upper endpoint
Step 5 of 5: Construct the 90 % confidence interval for the slope. Round your answers to three decimal places. Answer(How to Enter) 2 Points Keypad |囲Tables Lower endpoint: Upper endpoint:
Answer #1

Solution:

We are given the data on the number of bids an items on eBay received and the items selling price.

Part 3) We have to find the estimated variance of slope, that is: S2b1

S_{b1}=rac{S_{e}}{sqrt{SS_{xx}}}

SSE Se- n-2

SSE = SS_{yy}-rac{SS_{xy}^{2}}{SS_{xx}}

ry

SS_{xx}=sum x^{2} : : -: : left ( sum x imes sum x : : /: : n: : ight )

SS_{yy}=sum y^{2} : : -: : left ( sum y imes sum y : : /: : n: : ight )

Thus we need to make following table:

x: Price y: Number of Bids x^2 y^2 xy
29 2 841 4 58
32 3 1024 9 96
40 7 1600 49 280
45 8 2025 64 360
49 9 2401 81 441
sum x = 195 sum y = 29 r2 = 7891 y-= 207 sum xy =1235

Thus

ry

SSry 1235195 x 29/5 )

SSy1235 - 1131

SS_{xy}=104

SS_{xx}=sum x^{2} : : -: : left ( sum x imes sum x : : /: : n: : ight )

SS_{xx}=7891 : : -: : left ( 195 imes 195 : : /: : 5: : ight )

SS_{xx}=7891 : : -: : 7605

S S TT-286

SS_{yy}=sum y^{2} : : -: : left ( sum y imes sum y : : /: : n: : ight )

SS_{yy}=207 : : -: : left ( 29 imes 29 : : /: : 5: : ight )

SS- 207-168.2

SS_{yy}=38.8

Thus

SSE = SS_{yy}-rac{SS_{xy}^{2}}{SS_{xx}}

SSE = 38.8 -rac{104 ^{2}}{286 }

SSE = 38.8 -37.818182

SSE = 0.981818

SSE Se- n-2

S_{e}=sqrt{rac{0.981818 }{5-2}}

100 5 0.981818 Se-

Se-V0.327273

S_{e}=0.572078

S_{b1}=rac{S_{e}}{sqrt{SS_{xx}}}

S_{b1}=rac{0.572078 }{sqrt{286 }}

S_{b1}=rac{0.572078 }{16.911535 }

S_{b1}=0.03383

Thus estimated variance is given by:

S^{2}_{b1}=0.03383^{2}

S^{2}_{b1}=0.00114

S0.001

Part 4) 95% confidence interval for slope:

Formula:

(b_{1}-E: : ,: : b_{1}+E)

where

SSry SST bi

b_{1}=rac{104 }{286}

ь1 0.36364

and

E = t_{c} imes S_{b_{1}}

Thus we need to find t critical value for c = 95% confidence level.

Thus two tailed area = 1 - c = 1 - 0.95 = 0.05

and df = n - 2 = 5 - 2 = 3

Thus look in t table for df = 3 and two tail area = 0.05 and find t critical value.

Thus from t table , we get:

tc = 3.182

Thus

E = t_{c} imes S_{b_{1}}

E = 3.182 imes 0.03383

E = 0.10764

Thus

Lower end point = b_{1}-E=0.36364-0.10764=0.25600 = 0.256

Upper end point = b_{1}+E=0.36364+0.10764=0.47128 = 0.471

Lower end point = 0.256

Upper end point = 0.471

Part 5) 90% confidence interval for slope:

Formula:

(b_{1}-E: : ,: : b_{1}+E)

We need to find t critical value for c = 90% confidence level.

Thus two tailed area = 1 - c = 1 - 0.90 = 0.10

and df = n - 2 = 5 - 2 = 3

Thus look in t table for df = 3 and two tail area = 0.10 and find t critical value.

Thus from t table , we get:

tc = 2.353

Thus

E = t_{c} imes S_{b_{1}}

E- 2.353 x 0,03383

E = 0.07960

Thus

Lower end point = b_{1}-E=0.36364-0.07960 =0.28404 = 0.284

Upper end point = bi + E = 0.36364 0.07960 0.44323 = 0.443

Thus

Lower end point =0.284

Upper end point = 0.443